### v1

# R-PINN: Recovery-type a-posteriori estimator enhanced adaptive PINN

Abstract
In recent years, with the advancements in machine learning and neural networks, algorithms using physics-informed neural networks (PINNs) to solve PDEs have gained widespread applications. While these algorithms are well-suited for a wide range of equations, they often exhibit suboptimal performance when applied to equations with large local gradients, resulting in substantial localized errors. To address this issue, this paper proposes an adaptive PINN algorithm designed to improve accuracy in such cases. The core idea of the algorithm is to adaptively adjust the distribution of collocation points based on the recovery-type a-posterior error of the current numerical solution, enabling a better approximation of the true solution. This approach is inspired by the adaptive finite element method. By combining the recovery-type a-posteriori estimator, a gradient-recovery estimator commonly used in the adaptive finite element method (FEM) with PINNs, we introduce the Recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN) and compare its performance with a typical adaptive PINN algorithm, FI-PINN. Our results demonstrate that R-PINN achieves faster convergence with fewer adaptive points and significantly outperforms in the cases with multiple regions of large errors than FI-PINN. Notably, our method is a hybrid numerical approach for solving partial differential equations, integrating adaptive FEM with PINNs.
keywords:
Adaptive PINN , Recovery-type a-posteriori estimator , Machine learning , Adaptive FEM
1 Introduction
The Physics-informed neural networks (PINNs) algorithm is a data-driven approach for solving partial differential equations (PDEs) based on deep learning frameworks [1]. Due to its distinct methodology, which diverges significantly from traditional PDE numerical techniques, and by leveraging the superior capabilities of neural networks in handling large-scale data and optimization problems, PINNs have garnered considerable attention since their introduction and have become important tools for solving PDEs. These methods utilize automatic differentiation to compute the penalty of an equation at a set of collocation points within the domain of interest, using it as an objective loss function that is optimized through backpropagation, thereby approximating the solution of the PDE. Compared to classical PDE numerical methods such as the finite element method and finite difference method, one significant advantage of PINNs is that they are mesh-free, making them easier to implement and particularly well-suited for high-dimensional problems and irregular domains. Currently, PINNs have been applied to solve a range of PDE problems, demonstrating their strengths effectively, such as inverse problems in fluid dynamics, optics and metamaterials [2, 3], or some other forward problems [4, 5, 6].
Despite the notable advantages demonstrated by PINNs in PDEs, there remains substantial room for improvement in several aspects. While PINNs have exhibited high-precision solutions across a variety of scenarios, their applicability is still limited. Notably, significant errors can arise when addressing equations such as the Allen-Cahn equation or the Burgers’ equation. Enhancements to the PINNs methodology can be pursued from multiple aspects. For instance, since the total loss function in PINNs is a weighted sum of errors from both the domain and boundary conditions, adjusting the weights of these terms may improve performance across different problems [7, 8, 9, 10]. In cases involving extensive solution domains, decomposition techniques can be employed to localize the training process, thereby enhancing optimization outcomes [11, 12, 13]. Regarding the choice of loss functions, although Mean Squared Error (MSE) is commonly used, alternative forms of loss functions may give better training results in certain circumstances [14, 10, 15]. Furthermore, strategies such as batch training and the design of specialized network architectures to deal with Dirichlet or periodic boundary conditions offer a promising direction to enhance the performance of PINNs [8, 16, 17]. In this paper, we focus on the limitations of PINNs in handling solutions with sharp local variations. This challenge, arising from singularities or high-gradient regions, is typically addressed by introducing adaptive sampling strategies for PINNs. However, existing sampling methods primarily only rely on the values of the loss function at local collocation points, without incorporating a rigorous a-posteriori error estimator. To address this issue, we propose the recovery-type a-posteriori estimator-enhanced adaptive PINN (R-PINN), which integrates the recovery-type a-posteriori estimator from classical numerical methods into PINNs. Furthermore, we introduce an adaptive sampling method, the Recovery-type estimator adaptive distribution (R-AD), to systematically implement adaptive refinement.
The sampling methods for collocation points in PINNs are generally categorized into uniform and non-uniform sampling. In typical uniform sampling, points are either randomly distributed or placed at equispaced intervals, remaining fixed throughout the training process. However, this static sampling strategy has notable limitations, as it fails to capture the model’s fine local features effectively, which can lead to convergence issues in complex problems settings. To address these limitations, Lu et al. first proposed an adaptive non-uniform sampling method [18]. Since then, various adaptive sampling methods for collocation points have been developed [19, 20, 21, 22, 23, 24, 20]. Among these, the Failure-informed PINN (FI-PINN) by Gao et al. [20] is particularly notable, utilizing a truncated Gaussian model for adaptive sampling. Another significant contribution is by Wu et al. [19], who reviewed and compared a range of sampling techniques. Their work covers both uniform sampling methods, such as equispaced, random, and low-discrepancy sequences, and non-uniform methods, including residual-based adaptive refinement (RAR) and residual-based adaptive distribution (RAD), offering valuable guidance on selecting appropriate sampling strategies.
While there are various adaptive PINN methods, they generally share a common approach: sampling the uniform points as background collocation points and then adding non-uniform points as the adaptive collocation points based on an a-posteriori error estimator. Most of these methods utilize the residual-type a-posteriori estimator, which calculates the error as the loss of the neural network function. For adaptive finite element methods, two primary types of estimator are widely recognized: residual-type [25, 26] and recovery-type [27, 28, 29], with the latter based on gradient recovery. Although extensive research has been conducted on adaptive PINNs, there has been limited work on gradient recovery estimators. Therefore, we introduce an adaptive PINN enhanced by the recovery-type a-posteriori estimator.
The rest of this paper is organized as follows. Section 2 provides preliminary background, introducing the fundamental algorithm of Physics-informed neural networks (PINNs) and existing sampling strategies for both uniform and non-uniform purposes. In Section 3, we present the detailed algorithm of the proposed Recovery-type estimator-enhanced PINN (R-PINN). Section 4 presents our results using R-PINN on four representative singular problems, illustrating its performance and providing a comparative analysis. Finally, Section 5 summarizes our conclusions and outlines potential directions for future research.
2 Preliminaries
In this section, we will present the requisite preliminary knowledge. We will begin with an overview of PINNs algorithm in solving both forward and inverse PDEs problems. Following this, we will discuss some strategies for collocation points sampling employed in the implementation of PINNs, including both uniform and non-uniform sampling methodologies.
2.1 PINNs for solving forward and inverse PDEs
Analogous to the application of basis functions for approximating true solutions in traditional numerical methods for PDEs, the Physics-informed neural networks (PINNs) algorithm utilizes deep neural networks to approximate the exact solutions of equations, primarily employing fully connected feed-forward neural networks. We denote a fully connected feed-forward neural network with $L$ layers as $\mathcal{N}^{L}:\mathbb{R}^{N_{0}}\rightarrow\mathbb{R}^{N_{k}}$, where the $k$-th layer comprises $N_{k}$ neurons. The weight matrix and bias vector of the $k$-th layer (for $1\leq k\leq L$) can be represented as $\bm{W}^{k}\in\mathbb{R}^{N_{k}\times N_{k-1}}$ and $\bm{b}^{k}\in\mathbb{R}^{N_{k}}$, respectively. The input vector is denoted as $\bm{x}\in\mathbb{R}^{N_{0}}$, and the output vector of the $k$-th layer is expressed as $\mathcal{N}^{k}(\bm{x})$, especially, $\mathcal{N}^{0}(\bm{x})=\bm{x}$. The activation function $\Phi$, applied during the feed-forward transition from the $(k-1)$-th layer to the $k$-th layer (for $1\leq k\leq L$), introduces non-linearity into the network, thereby enhancing its learning capacity. Consequently, a feed-forward neural network can be mathematically defined as
$\mathcal{N}^{k}(\bm{x})=\Phi(\bm{W}^{k}\mathcal{N}^{k-1}(\bm{x})+\bm{b}^{k}),\quad 1\leq k\leq L.$ (2.1)
Letting $\tilde{\bm{\Theta}}=\{\bm{W}^{k},\bm{b}^{k}\}\in\mathcal{V}$ be all parameters of the neural network, including weights and biases, where $\mathcal{V}$ denotes the parameter space, we can finally express the neural network as
$\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}}),$
where $\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ emphasizes the dependence of the output layer $\mathcal{N}^{L}(\bm{x})$ on $\tilde{\bm{\Theta}}$. Initial parameters are typically generated according to a specified probability distribution.
In the context of solving PDEs with PINNs, we consider a PDE defined over the domain $\Omega\in\mathbb{R}^{d}$ with parameters $\bm{\lambda}$,
$f(\bm{x};u(\bm{x}))=f\left(\bm{x};\frac{\partial u}{\partial x_{1}},\cdots,\frac{\partial u}{\partial x_{d}};\frac{\partial^{2}u}{\partial x_{1}^{2}},\cdots,\frac{\partial^{2}u}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)=0,\quad\bm{x}=(x_{1},x_{2},\cdots,x_{d})\in\Omega,$ (2.2)
subject to the boundary condition:
$\mathcal{B}(u,\bm{x})=0\quad\text{on}\ \partial\Omega,$
where $u(\bm{x})$ represents the value of the true solution at point $\bm{x}$.
In solving the forward problem, where the parameter $\bm{\lambda}$ in the equation is known, PINNs approximate the true solution $u$ using a neural network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})$ [1]. During this process, the neural network is iteratively trained to update its parameters $\tilde{\bm{\Theta}}$ in order to minimize the loss function, defined as,
$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}}),$ (2.3)
where
$\begin{split}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{f}\rvert}\sum_{\bm{x}\in\mathcal{T}_{f}}\bigg{|}&f\left(\bm{x};\frac{\partial\hat{u}}{\partial x_{1}},\cdots,\frac{\partial\hat{u}}{\partial x_{d}};\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{1}},\cdots,\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)\bigg{|}^{2},\\ &\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{b}\rvert}\sum_{\bm{x}\in\mathcal{T}_{b}}\lvert\mathcal{B}(\hat{u},\bm{x})\rvert^{2},\end{split}$ (2.4)
and $w_{f}$ and $w_{b}$ are weighting factors. Here, $\mathcal{T}_{f}$ and $\mathcal{T}_{b}$ denote the sets of sampled points within the domain and on the boundary, respectively, which are commonly referred to as ”collocation points”. The combined set of collocation points is given by $\mathcal{T}=\mathcal{T}_{f}\cup\mathcal{T}_{b}$.
For certain problems, while the parameters $\bm{\lambda}$ in the governing equations remain unknown, the solutions are observable, thereby giving rise to inverse problems. To solve the inverse problem with PINNs, the loss function not only accounts for the deviations dictated by physical laws but also incorporates an additional term to capture the residual between observed solution $u(\bm{x})$ and predicted solution $\hat{u}(\bm{x})$ over a set of collocation points $\mathcal{T}_{i}$:
$\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda})=\frac{1}{|\mathcal{T}_{i}|}\sum_{\bm{x}\in\mathcal{T}_{i}}|\hat{u}(\bm{x})-u(\bm{x})|^{2}.$ (2.5)
As a result, the total loss function for the inverse problem can be expressed as:
$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}},\bm{\lambda})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}},\bm{\lambda})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}},\bm{\lambda})+w_{i}\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda}),$ (2.6)
with the weight of the additional term $w_{i}$. Then the unknown parameters $\bm{\lambda}$ can be trained with the neural network parameters $\tilde{\bm{\Theta}}$ by the optimization process.
Note that in equations (2.3) and (2.6), it is crucial to position the set of collocation points within the domain of interest. While simple sampling methods, such as uniform and random sampling, are available, they are generally less effective in capturing the solution’s local features. Consequently, in certain cases, employing more advanced non-uniform sampling techniques becomes particularly important.
2.2 Sampling methods for PINNs
Typically, the sampling strategy in PINNs involves selecting a set of collocation points by uniform sampling prior to training the neural network, with the set remaining fixed throughout the training process. Commonly employed uniform sampling methods include equispaced uniform sampling and uniformly random sampling, while other uniform sampling techniques are also available [1, 4, 30]. In Wu. et al.’s work [19], six uniform sampling methods were systematically reviewed, including equispaced uniform grid (Grid), uniformly random sampling (Random), Latin hypercube sampling (LHS) [31, 32], Halton sequence (Halton) [33], Hammersley sequence (Hammersley) [34], and Sobol sequence (Sobol) [35]. Their comparative experiments demonstrated that the Sobol method generally provides superior performance. The Sobel sequence uses a base of two to form successively finer uniform partitions of the unit interval and then reorder the coordinates in each dimension. We show the difference over three different methods in Fig. 1.
Fig. 1: $900$ uniformly collocation points generated on $[0,1]^{2}$ with three different strategies.
In many cases, accuracy of numerical solution can be enough through uniform sampling strategies. However, for certain numerical problems characterized by large local gradients and highly distinctive solution features, such methods may fail to adequately capture local variations. Several adaptive methods have been proposed to overcome these limitations in recent years, primarily through modifications to the sampling strategies for the collocation points in PINNs. These adaptive sampling approaches can be broadly categorized into two types [19]: the Residual-based adaptive refinement (RAR) [18] and the Residual-based adaptive distribution (RAD) [21]. RAR targets regions with large residuals, where new adaptive collocation points are introduced at regular intervals during the training process to reduce local errors (see Algorithm 1).
Algorithm 1 Residual-based adaptive refinement (RAR)
1:  Sample background collocation points $\mathcal{T}$.
2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
3:  while not reach the end do
4:     Sample a set of adaptive collocation points $\mathcal{S}_{0}$.
5:     Compute the PDE residuals of adaptive collocation points $\mathcal{S}_{0}$.
6:     $S\leftarrow m$ points with largest residuals from $\mathcal{S}_{0}$.
7:     $\mathcal{T}\leftarrow\mathcal{T}\cup\mathcal{S}$.
8:     Train the neural network for some epoches.
9:  end while
10:  Output the result.
While RAR only focuses on the area with large residuals and ignore the other place, RAD utilizes the probabilty density function $p(\bm{x})$ that has strong relationship with residual to resample all the collocation points, which can take the area both with large and small residual into care (see Algorithm 2). For example, we can take $p(\bm{x})$ proportional to the PDE residual $\epsilon(\bm{x})=|f(\bm{x};\hat{u}(\bm{x})|$,
[…]
More forms of $p(x)$ can be used here, such as it in paper[19],
$p(\bm{x})\propto\frac{\epsilon^{k}(\bm{x})}{\mathbb{E}[\epsilon^{k}(\bm{x})]}+c,$ (2.8)
where $k\geq 0,c\geq 0$ are two hyperparameters.
Algorithm 2 Residual-based adaptive distribution (RAD)
1:  Sample background collocation points $\mathcal{T}$.
2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
3:  while not reach the end do
4:     $\mathcal{T}\leftarrow$ resampled collocation points according to the probability density function $p(\bm{x})$.
5:     Train the neural network for some epoches.
6:  end while
7:  Output the result.
3 Recovery-type estimator enhanced PINN
In this section, we present an adaptive PINNs algorithm developed by our team, which utilizes the recovery-type a-posteriori error estimator [27, 28, 29], a widely used technique in adaptive finite element methods. The concept of adaptive finite element method, first introduced by I. Babuška et al. in 1978 [25], has since seen significant advancements and remains integral to practical scientific computations. This approach extends classical finite element methods by incorporating a-posteriori error estimation and automated mesh refinement to iteratively improve the accuracy of approximate solutions, which can achieve the target accuracy with reduced computational effort, offering distinct advantages over traditional finite element methods, particularly when addressing problems involving singular solutions.
The fundamental concept of the adaptive finite element method lies in the use of a-posteriori error estimator to assess the numerical solution error within each mesh element, thereby guiding the mesh refinement and coarsening process. This adaptive mesh strategy shares a notable similarity with adaptive PINNs, where the distribution of collocation points is adjusted based on the error distribution. In adaptive finite element methods, commonly employed a-posteriori error estimators can be categorized into two main types: the first is based on residual type [25, 36], where the error is estimated by computing the discrepancy between the solution and the residual of the governing equation; the second type, recovery type, proposed by the Zienkiewicz and Zhu [27, 28, 29], is a gradient reconstruction-based method that provides more accurate error estimates by reconstructing the solution gradients.
In existing adaptive PINNs algorithms, the equation operator is typically used as the default a-posteriori error estimator, which corresponds to the residual-based error estimator in adaptive finite element methods. Our proposed algorithm seeks to introduce the recovery-type error estimator into the adaptive PINNs framework, with the aim of enriching and extending the theoretical foundation of adaptive PINNs, ultimately providing a more comprehensive approach to error estimation and control.
3.1 Recovery-type a-posteriori estimator
In 1987, O.C. Zienkiewicz and J.Z. Zhu introduced a pioneering a-posteriori error estimator based on gradient recovery for finite element theory, commonly referred to as the Z-Z method [27] or the recovery-type method. This approach, notable for its computational simplicity, intuitive understanding, and straightforward implementation within finite element software, quickly gained widespread acceptance within the engineering community. Its rapid adoption represented a significant milestone, facilitating the transition from theoretical research to practical application in the domain of finite element a-posteriori error estimation.
In the context of a specific recovery-type a-posteriori error estimation framework, a typical problem in the field of elasticity is to be considered as an example, which is encapsulated within the following system of equations
$\bm{L}\bm{u}=\bm{S}^{T}\bm{DS}\bm{u}=\bm{f}\quad\text{in}\ \Omega,$ (3.1)
which is subject to suitable boundary conditions on the domain’s boundary $\partial\Omega$. Here, $\bm{u}$ represents the displacement vector, $\bm{L}$ is the differential operator intrinsic to the theory of elasticity, and the matrix operator $\bm{S}$ defines the strain tensor $\bm{\epsilon}$ by
$\bm{\epsilon}=\bm{S}\bm{u}.$ (3.2)
Furthermore, the elasticity matrix $\bm{D}$ is utilized to determine the stress tensor $\bm{\sigma}$, as delineated by the equation:
$\bm{\sigma}=\bm{D}\bm{\epsilon}.$ (3.3)
The finite element solution to the aforementioned elasticity problem is expressed in the following form
$\bm{u}_{h}=\bm{N}\bar{\bm{u}}_{h},$ (3.4)
where $\bar{\bm{u}}_{h}$ denotes the vector of nodal displacements, and $\bm{N}$ represents the set of approximate polynomial shape functions that interpolate the displacement field within the finite element framework Subsequently, the stress tensor $\bm{\sigma}_{h}$, is computed using the approximate displacement field $\bm{u}_{h}$ with
$\bm{\sigma}_{h}=\bm{DSu}_{h}.$ (3.5)
The recovery-type a-posteriori error estimator is characterized by focusing on estimating the numerical error by approximating the gradient recovery of the numerical solution $\bm{\sigma}^{*}$. Therefore, the gradient recovery technique plays a key role for the estimator. With the development of recovery-type a-posteriori error estimators, there have been many gradient recovery methods, the typical ones are the transformation method and nodal averaging method.
Fig. 2: The finite element method solution and the smoothened solution of stress for a one-dimensional problem.
  • 1.
    The transformation method
    For this method, it is assumed that the stress
    $\bm{\sigma}^{*}=\bm{N}\bar{\bm{\sigma}}^{*},$ (3.6)
    and the method seeks to minimize the discrepancy between the smooth stress $\bm{\sigma}^{*}$ and the stress of the finite element $\bm{\sigma}_{h}$ by enforcing the condition
[…]
    By combining equations (3.6) and (3.7), we can derive the following expression for $\bar{\bm{\sigma}}^{*}$
[…]
    This method is highly accurate but computationally expensive due to the need to solve the additional system of equations.
  • 2.
    The nodal averaging method
    In this simpler approach, the smooth stress $\bar{\sigma}^{*}$ is obtained by averaging the finite element stresses $\bm{\sigma}_{h,i}$ at the nodes
    $\bm{\sigma}^{*}=\frac{\left(\sum_{i=1}^{M}\bm{\sigma}_{h,i}\right)}{M}$ (3.9)
    where $M$ is the number of elements connected with the node $i$.
    This method is straightforward and computationally efficient, although it may not achieve the same level of accuracy as the transformation method. Nevertheless, it performs well in practical applications.
As for error estimation, two types of errors are usually considered. One is the error in $\bm{\sigma}$ which can be expressed as
$\lVert\bm{e}_{\sigma}\rVert=\lVert\bm{\sigma}-\bm{\sigma}_{h}\rVert.$ (3.10)
The other is the recovery-type estimator of the error is similarly defined in the appropriate norm as
$\lVert\tilde{\bm{e}}_{\sigma}\rVert=\lVert\bm{\sigma}^{*}-\bm{\sigma}_{h}\rVert,$ (3.11)
and its efficiency is gauged by the efficacy index defined as
$\theta=\frac{\lVert\tilde{\bm{e}}_{\sigma}\rVert}{\lVert\bm{e}_{\sigma}\rVert}.$ (3.12)
The confines of this efficacy index are elucidated through the subsequent theorem (as detailed in [29]):
$1-\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert}\leq\theta\leq 1+\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert},$ (3.13)
with the error in the reconstructed solution characterized as
$\lVert\bm{e}_{\sigma}^{*}\rVert=\lVert\bm{\sigma}-\bm{\sigma}^{*}\rVert.$ (3.14)
It is evident that the efficacy index approaches unity when the precision of the reconstructed solution significantly outperforms that of the original finite element solution.
Should the reconstructed solution exhibit superconvergence, that is,
$\lVert\bm{e}_{\sigma}^{*}\rVert=C^{*}h^{p+\alpha},$ (3.15)
where $\alpha$ is a positive constant and $h$ represents the element size, and
$\lVert\bm{e}_{\sigma}\rVert=Ch^{p},$ (3.16)
then the bounds for the efficacy index are
$1-\tilde{C}h^{\alpha}\leq\theta\leq 1+\tilde{C}h^{\alpha},$ (3.17)
with $\tilde{C}=C^{*}/C$, rendering the error estimator asymptotically precise. This estimation underscores the efficiency of the recovery-type a-posteriori estimator. Given that the error estimation revolves around stress $\bm{\sigma}$, which is contingent upon the gradients of the solution $\bm{u}$, this class of estimator may also be regarded as a form of gradient recovery.
3.2 Algorithm of Recovery-type estimator enhanced PINN
In this section, we introduce the algorithm of recovery-type estimator enhanced adaptive PINN (R-PINN). The success of the adaptive PINNs algorithm hinges on two crucial components: the posterior error estimation method and the adaptive sampling strategy. The posterior error estimation predicts the discrepancy between the numerical and true solutions, while the adaptive sampling strategy adjusts the distribution of collocation points based on this error, allowing for more focused and efficient training in areas that need it most. Section 3.1 provides a comprehensive explanation of the posterior error estimation method. Therefore, for our R-PINN, the key is how to combine PINNs with this kind of posterior error estimator to realize the adaptive process.
To solve the equation (2.2), we utilize a shape-regular triangulation $\mathcal{T}_{h}$ formed by a family of partitions of the domain $\Omega$, which allows us to effectively implement the recovery-type a-posteriori estimator. We define the operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ as the recovery-type estimator used in our algorithm, employing a nodal averaging method for gradient recovery (other gradient recovery methods can also be utilized). The approximated solution $\hat{u}_{\tilde{\Theta}}$ predicts the solution values at all nodes within the mesh $\mathcal{T}_{h}$, which then facilitates gradient recovery and error computation for each element $K\in\mathcal{T}_{h}$ through the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$.
Fig. 3: The flow chart of R-PINN.
Another crucial aspect is the selection of collocation points and the adaptive adjustment of their distribution. Similar to many adaptive PINNs algorithms, we start with $N_{1}$ background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ for pre-training. During the adaptive training phase, we introduce $N_{2}$ adaptive collocation points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ alongside the background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$, updating the adaptive collocation points after each training batch. For background collocation points, we utilize a Sobol sequence as described in section 2.2. For the adaptive collocation points, we propose a novel approach called the recovery-type estimator based adaptive distribution (RE-AD).
Unlike common sampling methods, the RE-AD technique is element-based, as the computed a-posteriori errors are distributed across elements. This means we only need to specify the number of adaptive collocation points to sample in each element, allowing for random sampling within each element. Therefore, it’s vital to allocate adaptive collocation points intelligently to target areas with significant errors without creating redundancy. To achieve this, we design an iterative process: we define a tolerance $\epsilon$ such that $\frac{N_{2}}{N_{e}}\leq\epsilon<1$, where $N_{e}$ is the total number of elements. Initially, we select $L_{1}=[\epsilon N_{e}]$ elements that exhibit the largest errors in $\mathcal{T}_{h}$, which we refer to as the sequence $\mathcal{S}_{1}$. Adaptive collocation points are sampled from $\mathcal{S}_{1}$ in proportion to the errors. For instance, for the $i$-th element in $\mathcal{S}_{1}$, the number of adaptive collocation points assigned is $[N_{2}\frac{e_{i}}{\sum_{j}e_{j}}]$. Since this process rounds down, some adaptive collocation points may remain unallocated. We then repeat this process by selecting $L_{2}=[\epsilon L_{1}]$ of the largest elements and allocating adaptive collocation points according to their errors, continuing until no points are redundant. The RE-AD sampling method differs from RAR but is more akin to RAD, as it maintains a fixed total number of adaptive collocation points while only adjusting their distribution during each iteration.
Algorithm 3 Recovery-type estimator based adaptive distribution (RE-AD)
1:  Set the number of redundant adaptive collocation points $N_{r}=N_{2}$.
2:  Denote $L_{0}=N_{e}$.
3:  $k\leftarrow 1$.
4:  while $N_{r}$ not 00 do
5:     Select $L_{k}=[\epsilon L_{k-1}]$ elements with the largest errors as $\mathcal{S}_{k}$.
6:     Designate adaptive collocation points to $\mathcal{S}_{k}$.
7:     Update the number of redundant adaptive collocation points $N_{r}$.
8:     $k\leftarrow k+1$.
9:  end while
10:  Output adaptive collocation points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$.
To introduce the specific procedures of the recovery-type estimator enhanced adaptive PINN (R-PINN), we utilize the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ along with the recovery-type estimator based adaptive distribution (RE-AD) sampling strategy. Initially, we use a set of background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ to train the neural network, aiming to obtain a rough approximation of the true solution, which serves as the pre-training step. Following the pre-training phase, an adaptive training iteration process begins. In each iteration, the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ is used to compute the errors of the approximated solution $\hat{u}_{\tilde{\Theta}}(\bm{x})$ on the mesh $\mathcal{T}_{h}$. Based on these errors, the RE-AD sampling method is then employed to strategically distribute additional adaptive collocation points. These adaptive collocation points are combined with the background collocation points to retrain the neural network. This iterative process continues until the accuracy improvements become minimal, indicating that the network has converged to a satisfactory approximation of the solution. Since the number of collocation points can always stay constant, the algorithm satisfies the characteristics of self-adaptation while ensuring accurate solution.
Algorithm 4 Recovery-type estimator enhanced PINN (R-PINN)
0:  The number of adaptive samples and background samples $N_{a}$ and $N_{b}$, the maximum steps number $M$
1:  Generate $N_{1}$ background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ with Sobol sequence.
2:  Pre-train the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$.
3:  Set the number of iteration $k\leftarrow 1$ and give the max number of iterations $M$.
4:  while $k\leq M$ do
5:     Generate $N_{2}$ adaptive samples $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ with RE-AD method.
6:     Retrain the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with the dataset $\mathcal{D}=\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}\cup\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$.
7:  end while
[…]
4 Numerical experiments
In this section, we present numerical experiments demonstrating the effectiveness of our proposed algorithm. Our evaluation compares three different kinds of PINNs: the Recovery-type estimator enhanced PINN (R-PINN), Failure-informed PINN (FI-PINN) [20], and standard PINN as a baseline.
4.1 Experimental setup
To ensure a fair comparison, we maintain identical parameters and training processes across all methods for the same number of iterations, varying only the number and distribution of adaptive collocation points. The recovery-type a-posteriori estimation is implemented using FEALPy [37].
All experiments employ a fully-connected neural network with $7$ hidden layers, each containing 20 neurons with tanh for the activation function unless otherwise specified. The training process comprises two phases: an initial pre-training phase utilizing the L-BFGS optimizer with a learning rate of $0.1$, conducted for a problem-dependent number of epochs, followed by an adaptive training phase consisting of $M$ adaptive iterations. Each iteration is constrained to $5000$ epochs while employing the same optimization method.
For collocation points, we use $N_{1}=2000$ background collocation points and $N_{2}=100$ adaptive collocation points for R-PINN unless specified, while FI-PINN requires $N_{2}=500$ adaptive collocation points to achieve convergence unless specified. For the standard PINN, only $N_{1}=2000$ background collocation points are used during training, without incorporating additional adaptive collocation points. However, the training process remains the same as that of the two adaptive PINNs, with each adaptive iteration also consisting of 5000 epochs. All methods utilize $200$ boundary collocation points. Notably, our approach achieves comparable or superior accuracy despite using fewer total epochs and adaptive collocation points than related studies. The larger $N_{2}$ value for FI-PINN is necessary as it fails to achieve adaptivity with fewer points. Furthermore, while our method maintains a fixed number of adaptive collocation points, FI-PINN accumulates points throughout the adaptive iteration process. To evaluate the accuracy of our method, we employ two error metrics: the relative $L^{2}$ error,
$\frac{\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{2}}{\lVert u\rVert_{2}},$ (4.1)
and the $L^{\infty}$ error, defined as $\lVert\hat{u}{\tilde{\Theta}}-u\rVert_{\infty}$. Here, $\hat{u}_{\tilde{\Theta}}$ denotes the numerical solution obtained by our method, and $u$ represents the exact solution.
4.2 Poisson equation
Consider the following Poisson equation:
$\begin{split}-\Delta u(x,y)=&f(x,y),\quad(x,y)\ \text{in}\ \Omega,\\ u(x,y)=&g(x,y),\quad(x,y)\ \text{on}\ \partial\Omega,\end{split}$ (4.2)
where the domain $\Omega$ is $=[0,1]\times[0,1]$, and the true solution is
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}.$ (4.3)
The solution exhibits a sharp peak near the point $(0.5,0.5)$, which is challenging for standard PINN to capture accurately. In this experiment, we pre-train the network for $10000$ iterations and set the maximum number of adaptive iterations to $M=5$.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $3.386$ $0.0990$ $100$ $2.717$ $0.6541$ $500$ $0.809$ $0.6270$
$2$ $0.732$ $0.0282$ $100$ $1.514$ $0.4593$ $1000$ $0.689$ $0.6180$
$3$ $0.125$ $0.0260$ $100$ $1.157$ $0.3095$ $1500$ $0.650$ $0.5683$
$4$ $0.285$ $0.0088$ $100$ $0.294$ $0.0686$ $2000$ $0.879$ $0.5689$
$5$ $0.139$ $0.0086$ $100$ $0.068$ $0.0335$ $2500$ $0.682$ $0.5570$
Table 1: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving Poisson equation.
R-PINN
(a) *
FI-PINN
(b) *
Fig. 4: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods.
Fig. 4 illustrates the distribution of adaptive collocation points by two adaptive methods in the first two adaptive iterations. It is evident that all adaptive collocation points of our model converge to the neighborhood of the point $(0.5,0.5)$, while adaptive collocation points of FI-PINN are more dispersed. This shows the strong effectiveness of our algorithm in handling problems with significant local variations. Table 1 compares the relative $L^{2}$ error and $L^{\infty}$ error among R-PINN, FI-PINN and standard PINN during the adaptive iterations where $N_{\text{iter}}$ represents the number of adaptive iterations. The relative $L^{2}$ error for R-PINN reaches $0.139$, and the $L^{\infty}$ error reaches $0.0086$. In contrast, the errors obtained by FI-PINN decrease slowly at the beginning, but drop fast later, reaching the relative $L^{2}$ error of $0.068$ and $L^{\infty}$ error of $0.0335$. It can be observed that R-PINN can converge faster than FI-PINN. While, R-PINN can reach a smaller relative $L^{2}$ error in this case with the exponentially accumulated adaptive collocation points, which is one of its advantage. As for the normal PINN, the errors obtained show little to no decrease, indicating that the normal PINN fails to converge effectively. Similar results are observed in Fig. 6, where the lowest relative $L^{2}$ error for our algorithm is below $1\times 10^{-1}$. In the left panel of Fig. 5, we show the absolute errors throughout the entire domain, clearly demonstrating that the errors are much smaller with R-PINN. The middle panel of Fig. 5 illustrates the cross-section at $y=0.5$. The predicted curve using R-PINN fits the true solution very well, while the solution by FI-PINN has a little obvious error at the foot of the peak. The final solutions with three methods are plotted in the right of Fig. 5.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 5: The prediction err (left), the cross section at $y=0.5$ (middle) and the predicted solution of three methods solving Poisson equation.
Fig. 6: The relative $L^{2}$ errors of three methods in the optimization process of solving Poison equation.
4.3 Burgers’ equation
We consider the following Burgers’ equation:
$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\frac{0.01}{\pi}\frac{\partial^{2}u}{\partial x^{2}},\\ &u(x,0)=-\sin(\pi x),\\ &u(t,-1)=u(t,1)=0\end{split}$ (4.4)
where the domain is $(x,t)\in[-1,1]\times[0,1]$. The solution to this Burgers’ equation exhibits a sharp mutation along $x=0$, making it prone to errors and offsets. For this test, we pre-train the network for $5000$ iterations and use $M=4$ steps of adaptive processes.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.0086$ $0.0735$ $100$ $0.2278$ $1.8142$ $200$ $0.0136$ $0.1583$
$2$ $0.0107$ $0.0964$ $100$ $0.0131$ $0.1263$ $400$ $0.0144$ $0.1685$
$3$ $0.0071$ $0.0673$ $100$ $0.0064$ $0.0576$ $600$ $0.0174$ $0.2099$
$4$ $0.0059$ $0.0439$ $100$ $0.0089$ $0.0846$ $800$ $0.0197$ $0.2418$
Table 2: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving Burgers’ equation.
R-PINN
(a) *
FI-PINN
(b) *
Fig. 7: The distribution of adaptive collocation points in the first $2$ adaptive iterations with two adaptive PINN methods solving Burgers’ equation.
The adaptive collocation points generated during the first two adaptive training processes by two adaptive PINN method are illustrated in Fig. 7. We observe that the adaptive collocation points successfully cluster around $x=0$ with both methods, where the error is largest (as seen in Fig. 8). This clustering significantly improves the local error. The exact relative $L^{2}$ error and $L^{\infty}$ error are listed in Table 2. As shown, the final relative $L^{2}$ error and $L^{\infty}$ error for R-PINN are $0.0059$ and $0.0439$, respectively, while those for FI-PINN are $0.0089$ and $0.0846$. The prediction err of standard PINN keeps at the save level during the training, where relative $L^{2}$ error and $L^{\infty}$ reach $0.0197$ and $0.2418$ in the final. Similar to the experiment of Poisson equation, it takes a long iterations for FI-PINN before the error jumps quickly, while R-PINN can reduce the errors at once. The errors obtained by R-PINN and FI-PINN are one order of magnitude lower than those obtained by the standard PINN. Such behaviour can also be observed in the Fig. 10, where the relative $L^{2}$ errors are illustrated specifically, showing that the lowest error can reach below $1\times 10^{-2}$. We also show the absolute errors across the entire domain and the cross-section at $t=1$ in Fig. 8. The predicted solutions of two adaptive methods are illustrated in Fig. 9.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 8: The prediction err (left) and the cross section at $t=1$ (right) of three methods solving Burgers’ equation.
Fig. 9: The predicted solution (R-PINN at the left and FI-PINN at the right).
Fig. 10: The relative $L^{2}$ errors of three methods in the optimization process of solving Burgers’ equation.
4.4 Two-dimensional problem with two peaks
Considering the following two-dimensional problem
$\begin{split}-\nabla\cdot[u(x,y)\nabla(x^{2}+y^{2})]+\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in\Omega,\\ u(x,y)=b(x,y),\quad(x,y)\in\partial\Omega,\end{split}$ (4.5)
where $\Omega=[-1,1]\times[-1,1]$. The true solution is
$u(x,y)=e^{-1000[(x-0.5)^{2}+(y-0.5)^{2}]}+e^{-1000[(x+0.5)^{2}+(y+0.5)^{2}]},$ (4.6)
which has two peaks at $(-0.5,-0.5)$ and $(0.5,0.5)$. This test mainly can show the ability of R-PINN to deal with multiple peaks. For this experiment, we pre-train our network for the period of $10000$ iterations and set the maximum number of adaptive iterations as $M=4$. Since two peaks problem is challenging for neural network to approximate, we use $5$ hidden layers with $64$ neurons in each layer to enhance its capability. Moreover, we use $N_{2}=150$ adaptive collocation points for R-PINN, and $N_{2}=750$ added adaptive collocation points for FI-PINN.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.6778$ $0.1228$ $150$ $0.4372$ $0.4845$ $750$ $0.7912$ $0.6651$
$2$ $0.1202$ $0.0198$ $150$ $0.5220$ $0.3429$ $1500$ $0.8020$ $0.6475$
$3$ $0.1080$ $0.0259$ $150$ $0.4773$ $0.2233$ $2250$ $0.8012$ $0.6483$
$4$ $0.0840$ $0.0127$ $150$ $0.3969$ $0.2263$ $3000$ $0.8342$ $0.6625$
Table 3: The relative $L^{2}$ error and $L^{\infty}$ error by two methods solving two-dimensional problem with two peaks.
R-PINN
(a) *
FI-PINN
(b) *
Fig. 11: The distribution of adaptive collocation points in the first two adaptive iterations of two adaptive methods solving the two-dimensional problem with two peaks.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 12: The predicted err (left), the cross section at $y=-0.5$ (middle) and the cross section at $y=0.5$ (right) of three methods solving the two-dimensional problem with two peaks.
Fig. 13: The predicted solution (R-PINN at the left, FI-PINN at the middle and standard PINN at the right) of the two-dimensional problem with two peaks.
Fig. 14: The relative $L^{2}$ errors of three methods in the optimization process of solving the two-dimensional problem with two peaks.
We present the distribution of adaptive collocation points of two adaptive methods across the first two adaptive iterations in Fig. 11. It is evident that these points converge around the vicinities of $(0.5,0.5)$ and $(-0.5,-0.5)$ with R-PINN, demonstrating the capability of our algorithm to effectively handle problems characterized by multiple regions of abrupt changes. However, the distribution of adaptive collocation points with FI-PINN seems not so effective to deal with the two peaks. The reason of it might be a truncated Gaussian distribution is used, so that problems of more than one regions with large error are difficult to handle. This problem is also reflected in the training process. Table 3 compares the relative $L^{2}$ and $L^{\infty}$ errors among R-PINN, FI-PINN and standard PINN during the adaptive iterations. For R-PINN, the relative $L^{2}$ error reduces to $0.0840$, and the $L^{\infty}$ error decreases to $0.0127$. In contrast, the relative $L^{2}$ error associated with FI-PINN can only reach $0.3969$ and the $L^{\infty}$ can get to $0.2263$, which is just a little bit lower than those with standard PINN. Similar trends are observed in Fig. 14, where the lowest relative $L^{2}$ error for R-PINN falls below $1\times 10^{-1}$.
Fig. 12 provides additional insights. The left panel illustrates the absolute errors across the entire domain, revealing significantly smaller errors for R-PINN. The middle and right panels display cross-sectional views at $y=-0.5$ and $y=0.5$, respectively, showing that the predictions from R-PINN align closely with the exact solution, whereas there are large errors near two peaks with FI-PINN and the standard PINN can only approximate the features. Finally, the numerical solutions obtained by the three methods are visualized in Fig. 13.
4.5 Wave equation
For this test, we consider a wave equation as follow,
$\begin{split}&\frac{\partial^{2}u}{\partial t^{2}}-3\frac{\partial^{2}u}{\partial x^{2}}=0,\quad(t,x)\in[0,6]\times[-5,5],\\ &u(0,x)=\frac{1}{\cosh(2x)}-\frac{0.5}{\cosh(2(x-10))}-\frac{0.5}{\cosh(2(x+10))},\\ &\frac{\partial u}{\partial t}(0,x)=0,\\ &u(t,-5)=u(t,5)=0,\end{split}$ (4.7)
where the true solution is
$u(t,x)=\frac{0.5}{\cosh(2(x-\sqrt{3}t))}-\frac{0.5}{\cosh(2(x-10+\sqrt{3}t))}+\frac{0.5}{\cosh(2(x+\sqrt{3}t))}-\frac{0.5}{\cosh(2(x+10-\sqrt{3}t))}.$ (4.8)
Fig. 15: The plot of the exact solution.
R-PINN
(a) *
FI-PINN
(b) *
standard PINN
(c) *
Fig. 16: The predicted solution (left), predicted error (middle) and the distribution of adaptive collocation points (right) of three methods solving the wave equation.
Fig. 17: The relative $L^{2}$ errors of three methods in the optimization process of solving the Poison equation the wave equation.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $L^{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.0572$ $0.0761$ $100$ $0.1110$ $0.1590$ $500$ $0.3448$ $0.4554$
$2$ $0.0076$ $0.0098$ $100$ $0.0317$ $0.0491$ $1000$ $0.3158$ $0.4214$
Table 4: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving the wave equation.
In this test, the R-PINN method quickly reaches its performance limit. Therefore, we adopt a single adaptive iteration process ($M=2$) and pre-train the network for 5000 iterations. We show the relative $L^{2}$ errors and $L^{\infty}$ errors in the Table 4. The relative $L^{2}$ error and $L^{\infty}$ error can reach $0.0076$ and $0.0098$ with R-PINN, which are obviously lower than the other two methods. The detailed decreasing process is convenient to be seen in Fig. 17 The exact solution is depicted in Fig. 15. The left panel of Fig. 16 compares the predicted solutions of the three methods. The middle panel of Fig. 16 illustrates the absolute errors across the domain, showing that the errors associated with R-PINN are smaller than those with FI-PINN and standard PINN. Finally, the right panel of Fig. 16 displays the adaptive collocation points for two adaptive methods and random points for the standard PINN. From the right panel, we can see that the same problem occurs for FI-PINN when dealing with multiple large errors region, so it can’t handle well.
5 Conclusion
In this paper, we present recovery-type a-posteriori estimator enhanced adaptive PINN that combines the PINN with a recovery type estimator which is popularly used in the adaptive finite element method. Our innovation lies in the use of the recovery-type a-posteriori error estimator, which is different from the residual of PINNs, and the design of a tailored sampling method R-AD. We test the performance of R-PINN in several PDE problems with sharp or singular solutions, which normal PINN struggles to handle effectively. The results demonstrate that R-PINN achieves significant improvements over normal PINN. More studies can be done to analyze the performance of R-PINN for more complex problems and we shall focus on and analyze these issues in our future work.